### v1

@@ 4.2 Convergence
$\Phi_{t}\coloneqq(\Sigma_{x,t}^{-})^{-1}+C^{\top}\hat{\Sigma}_{v}^{-1}C-\left(I+X_{t}\Delta_{t}\right)^{-1}X_{t},$ (21)
and
$X_{t}:=(\Sigma_{x,t}^{-})^{-1}+C^{\top}(\Sigma_{v,t}^{*})^{-1}C.$
Assumption˜4 is a structural compatibility condition ensuring that the perturbation in the process-noise covariance lies in the range of the dynamics matrix $A$. While the DR Riccati iteration often converges numerically even when this condition is mildly violated, our formal convergence and uniqueness guarantees rely on it.
@@
Assumption˜4 holds whenever $\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$. In this case, a canonical choice of $\Delta_{t}$ is the minimum-Frobenius-norm solution
$\Delta_{t}:=A^{+}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})(A^{+})^{\top}.$ (22)
This matrix $\Delta_{t}$ is positive semidefinite, satisfies $A\Delta_{t}A^{\top}=\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}$, and obeys the bound
$\|\Delta_{t}\|_{2}\leq\|A^{+}\|_{2}^{2}\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}.$
In practice, verifying $\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$ requires knowledge of the time-varying covariance $\Sigma_{w,t}^{*}$. A convenient sufficient condition that can be checked a priori uses the spectral bound $\overline{\lambda}_{w}$ from Corollary˜2: if $\mathrm{range}(\overline{\lambda}_{w}I-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$, then Assumption˜4 holds for all $t$, since
$0\preceq\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\preceq\overline{\lambda}_{w}I-\hat{\Sigma}_{w}\implies\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(\overline{\lambda}_{w}I-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A).$
The form in (20) is structurally analogous to the risk-sensitive Riccati equation [4, 34]. Levy et al. [35] showed that, under Assumption˜3, the $N$-fold composition of the risk-sensitive Riccati mapping is a contraction in a suitable Riemannian metric for $N\geq n_{x}$, while Zorzi [36] extended this result to robust filters under the Thompson part metric. These results motivate the use of contraction-based arguments for analyzing the DR Riccati mapping.
An alternative interpretation of (20) arises by embedding the DR filter in a Krein space [37]. Augmenting the nominal system with a synthetic uncertainty $u_{t}$ and enforcing the constraint $0=\hat{x}_{t}+u_{t}$ yields an indefinite inner product on $u_{t}$ weighted by $\Phi_{t}^{-1}$. This perspective links distributional robustness to risk-sensitive filtering and provides a natural foundation for contraction-based analyses.
To formalize this idea, fix $N\in\mathbb{N}$ and consider the downsampled state $\hat{x}_{t}^{d}=\hat{x}_{tN}$. Downsampling lifts the time-varying recursion to an $N$-block time-invariant map, yielding block-diagonal perturbations and enabling a contraction argument similar to [37]. This induces an equivalent block model with matrices $(\mathcal{R}_{N},\mathcal{O}_{N},\mathcal{L}_{N},\mathcal{H}_{N})$ and a block-diagonal perturbation
$\bar{\Phi}_{N,t}=\mathrm{blkdiag}(\Phi_{tN+N-2},\ldots,\Phi_{tN-1}).$
The details are provided in Appendix C.1. Letting $\Sigma_{x,t}^{-,d}\coloneq\Sigma_{x,tN}^{-}$ denote the downsampled prior covariance matrix, the resulting Riccati recursion becomes
$\Sigma_{x,t+1}^{-,d}=r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d}),$ (23)
where $r_{\mathbb{D},t}^{d}$ denotes the $N$-fold composition of $r_{\mathbb{D}}$. If $r_{\mathbb{D},t}^{d}$ is a contraction for sufficiently large $t$, then so is $r_{\mathbb{D}}$. By Banach’s fixed-point theorem [38], this ensures the existence and uniqueness of a fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and hence of a steady-state gain $K_{\infty}$.
The following proposition provides a tractable sufficient condition for ensuring contraction.
Proposition 2.
Suppose Assumptions˜1, 2, 3 and 4 hold. Define
$\tilde{\phi}_{N}\;=\;\frac{1}{\lambda_{\max}\left(\mathcal{L}_{N}\bigl(I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N}\bigr)^{-1}\mathcal{L}_{N}^{\top}\right)}.$
Then, for some $N\geq n_{x}$, there exists $\phi_{N}\in(0,\tilde{\phi}_{N})$ such that if
$0\preceq\bar{\Phi}_{N,t}\preceq\phi_{N}I_{Nn_{x}},\qquad\forall t\geq q,$ (24)
for some $q>0$, the mapping $r_{\mathbb{D},t}^{d}$ is a contraction with respect to the Thompson part metric [39] on the cone of positive definite matrices for all $t\geq q$, and hence $r_{\mathbb{D}}$ is also a contraction in the same metric.
The proof is provided in Appendix C.2. A practical procedure for determining a suitable $\phi_{N}$ is described in [36]. One begins with $\phi_{N}=\tilde{\phi}_{N}$ and checks whether the associated Gramian $\Omega_{\phi_{N}I_{Nn_{x}}}$ is positive definite. If not, $\phi_{N}$ is iteratively reduced until $\Omega_{\phi_{N}I_{Nn_{x}}}$ becomes positive semidefinite but singular. At this threshold, if $\bar{\Phi}_{N,t}\prec\phi_{N}I_{Nn_{x}}$ for all $t\geq\tilde{q}$, then both Gramians $\Omega_{\bar{\Phi}_{N,t}}$ and $W_{\bar{\Phi}_{N,t}}$ are positive definite for $t\geq\tilde{q}$, which guarantees that the downsampled DR Riccati mapping (23) is a contraction.
Proposition˜2 establishes convergence provided that $\bar{\Phi}_{N,t}$ remains uniformly bounded by $\phi_{N}I_{Nn_{x}}$. In practice, however, $\bar{\Phi}_{N,t}$ depends on the sequence of SDP solutions in (7) and is therefore difficult to verify a priori. This motivates the derivation of explicit, tractable sufficient conditions on the ambiguity radii $(\theta_{v},\theta_{w})$ that guarantee the bound $\bar{\Phi}_{N,t}\preceq\phi_{N}I_{Nn_{x}}$ uniformly across time. Here, our noise-centric modeling becomes essential, as it leads to closed-form, computationally tractable expressions for these radii.
@@
Suppose Assumptions˜3, 1, 2 and 4 hold. Let $\phi_{v},\phi_{x}\geq 0$ with $\phi_{v}+\phi_{x}\leq\phi_{N}$, where $\phi_{N}$ is defined in Proposition˜2. Then, the following properties hold.
  1. (ii)
    (Measurement uncertainty bound) For any $\phi_{v}\geq 0$, define
    $\delta_{v}^{\max}(\phi_{v})\coloneqq\frac{\phi_{v}(\lambda_{\min}(\hat{\Sigma}_{v}))^{2}}{\|C\|_{2}^{2}+\phi_{v}\lambda_{\min}(\hat{\Sigma}_{v})}<\lambda_{\min}(\hat{\Sigma}_{v}).$ (25)
    Then, any $\theta_{v}$ satisfying
    $0\leq\theta_{v}\leq\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})+\delta_{v}^{\max}(\phi_{v})}-\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})}$ (26)
    guarantees that, for all $t\geq q$,
    $C^{\top}\!\bigl(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\bigr)C\;\preceq\;\phi_{v}I.$ (27)
  2. (iiii)
    (Process uncertainty bound) Define
    $\delta_{x}^{\max}\;\coloneqq\;\frac{1}{\lambda_{\min}(\hat{\Sigma}_{w})}+\frac{\|C\|_{2}^{2}}{\lambda_{\min}(\hat{\Sigma}_{v})}.$ (28)
    If $0\leq\phi_{x}<\delta_{x}^{\max}$, then any $\theta_{w}$ satisfying
    $0\leq\theta_{w}\leq\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})+\frac{\phi_{x}}{\|A^{+}\|_{2}^{2}\,\delta_{x}^{\max}(\delta_{x}^{\max}-\phi_{x})}}-\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}$ (29)
    guarantees that, for all $t\geq q$,
    $X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\preceq\phi_{x}I.$ (30)
Moreover, any pair $(\theta_{v},\theta_{w})$ satisfying (26) and (29) ensures that the DR Riccati mapping (19) admits a unique fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and the corresponding DR Kalman gain converges: $K_{t}\to K_{\infty}$.
The proof of this theorem is provided in Appendix C.4. The bounds (26) and (29) yield tractable, closed-form radii $(\theta_{v},\theta_{w})$ that certify convergence, directly in terms of nominal covariances and operator norms. This explicit tuning rule represents a practical consequence of our noise-centric ambiguity design.
@@
The radius $\theta_{x_{0}}$ is arbitrary and affects only the transient phase. The sufficient conditions (26) and (29) depend only on $(\theta_{v},\theta_{w})$ and the nominal noise covariances. In particular, for any $\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{++}$, the first prediction satisfies $\Sigma_{x,1}^{-}\succeq\Sigma_{w,0}^{\star}\succeq\hat{\Sigma}_{w}$. Thereafter, the contraction bounds are independent of the prior. Hence, if  (26) and (29) hold uniformly, one may take $q=0$ in Theorem˜3; otherwise, convergence is still guaranteed after a finite transient $q\geq 1$.
An immediate implication of Theorem˜3 is that the steady-state solution can be computed directly through the single stationary SDP (13) offline, rather than via online Riccati iterations.
Corollary 3.
Suppose that Assumptions˜1, 2, 3 and 4 hold. Then, the steady-state covariance matrices $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$, together with the least-favorable noise covariances $\Sigma_{w,\infty}^{*}$ and $\Sigma_{v,\infty}^{*}$, solve the stationary SDP (13). Consequently, $K_{\infty}^{*}=K_{\infty}=\lim_{t\to\infty}K_{t}$.
Under Assumption˜3, the Riccati recursions of the standard KFs in Theorem˜2 converge. When the DRKF covariances also converge, we can take limits in the matrix inequality of Theorem˜2 and, using that limits preserve positive semidefiniteness, obtain steady-state upper and lower bounds on the DRKF covariance, stated in the following corollary.
Corollary 4.
Suppose that Assumptions˜1, 2, 3 and 4 hold and the DRKF converges. Then, the steady-state prior and posterior error covariances $\Sigma_{x,\infty}^{-}$ and $\Sigma_{x,\infty}$ are bounded by the limits of the lower- and upper-bound Kalman recursions described in Theorem˜2:
$\underline{\Sigma}_{x,\infty}^{-}$ $\preceq\Sigma_{x,\infty}^{-}\preceq\overline{\Sigma}_{x,\infty}^{-}$
$\underline{\Sigma}_{x,\infty}$ $\preceq\Sigma_{x,\infty}\preceq\overline{\Sigma}_{x,\infty},$
where $\underline{\Sigma}_{x,\infty}^{-}\coloneqq\lim_{t\to\infty}\underline{\Sigma}_{x,t}^{-}$ and $\overline{\Sigma}_{x,\infty}^{-}\coloneqq\lim_{t\to\infty}\overline{\Sigma}_{x,t}^{-}$ denote the steady-state bounds for the prior covariances, and the posterior bounds $\underline{\Sigma}_{x,\infty}$ and $\overline{\Sigma}_{x,\infty}$ are defined analogously.
Finally, the DR Riccati recursion (18) reveals a structural connection to existing robust filtering paradigms. In particular, it resembles the risk-sensitive Riccati equation [4, 34]. This correspondence implies that the time-varying DRKF can be interpreted as a risk-sensitive filter with an effective risk parameter $\theta$ and a time-varying weighting matrix $Q_{t}\in\mathbb{S}^{n_{x}}_{+}$ satisfying $\theta Q_{t}^{\top}Q_{t}=\Phi_{t-1}$. At steady state, DRKF reduces to a risk-sensitive filter with a constant weighting matrix $Q$ chosen such that $\theta Q^{\top}Q=\Phi_{\infty}$, where $\Phi_{\infty}$ denotes the limiting perturbation.
4.3 Stability
A fundamental property of the classical KF is the asymptotic stability of its estimation error dynamics under standard controllability and observability assumptions. Specifically, the closed-loop error matrix (IKC)A(I-KC)A is Schur, ensuring boundedness of the estimation error covariance and decay of the error mean in the absence of bias.
In the DR setting, however, the estimator competes against the least-favorable noise distributions within the ambiguity set. A natural question is therefore whether the DRKF preserves this stability property. The following result shows that, despite the adversarial nature of the noise, the DRKF error dynamics remain asymptotically stable in the same sense as in the classical case.
Theorem 4.
Suppose Assumptions˜3, 1, 2 and 4 hold, and the ambiguity radii $(\theta_{v},\theta_{w})$ satisfy the conditions of Theorem˜3. Let $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$ and $(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*})$ denote the steady-state DRKF covariances, and let $K_{\infty}$ be the corresponding steady-state DR gain. Define the closed-loop error matrix $F_{\infty}\coloneq(I-K_{\infty}C)A$. Then, $F_{\infty}$ is Schur stable, and the posterior estimation error $e_{t}:=x_{t}-\bar{x}_{t}$ evolves according to
[…]
Consequently, under the least-favorable noise distributions,
$\lim_{t\to\infty}\mathbb{E}\|e_{t}\|^{2}\;=\;\operatorname{Tr}[\Sigma_{x,\infty}^{*}]\quad\text{and}\quad\sup_{t\geq 0}\,\mathbb{E}\|e_{t}\|^{2}\;<\;\infty.$
The proof is provided in Appendix D.1.
While Theorem˜4 guarantees mean-square stability, it does not by itself characterize the estimation-error mean in the presence of bias. When the least-favorable distributions (or the true noise processes) have nonzero means, the filter tracks a steady bias determined by the mean mismatch. The following corollary characterizes this bias.
Corollary 5.
Under the assumptions of Theorem˜4, let $m_{t}:=\mathbb{E}[e_{t}]=\mathbb{E}[x_{t}-\bar{x}_{t}]$ denote the posterior error mean under the least-favorable distributions. Let $\mathbb{E}[w_{t}]=\mu_{w}$ and $\mathbb{E}[v_{t}]=\mu_{v}$. Then,
[…]
and hence
[…]
In particular, if the noise means used by the filter coincide with the true means ($\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$), then $\mathbb{E}[e_{t}]\to 0$.
The proof of this corollary is provided in Appendix D.2. Corollary˜5 shows that our DRKF mirrors the classical KF in its treatment of bias: if the true noise means coincide with those used by the filter ($\mu_{w}=\hat{w}$ and $\mu_{v}=\hat{v}$), then the estimation-error mean vanishes asymptotically. If instead either mean is misspecified ($\mu_{w}\neq\hat{w}$ or $\mu_{v}\neq\hat{v}$), the filter remains stable and the bias converges to the constant vector in the right-hand side above, which depends only on $F_{\infty}$, $K_{\infty}$, and the mean mismatch.
4.4 Optimality
A key property of the classical KF is that it is the unique linear, causal estimator minimizing the MSE when the Gaussian noise distributions are known exactly. In the DR setting, however, the true noise covariances are unknown and only assumed to lie within Wasserstein ambiguity sets. In this case, it is natural to consider minimax optimality: among all causal estimators, the filter should minimize the worst-case one-step MSE over all distributions in the ambiguity sets.
The following theorem establishes that DRKF satisfies this minimax optimality in an asymptotic sense.
Theorem 5.
Suppose Assumptions˜1, 2, 3 and 4 hold. Let $(\theta_{v},\theta_{w})$ satisfy the conditions of Theorem˜3, and let $\Sigma_{x,\infty}^{-,*}\in\mathbb{S}^{n_{x}}_{++}$ denote the unique fixed point of the DR Riccati map $r_{\mathbb{D}}$. Let $\Sigma_{x,\infty}^{*}$ be the corresponding steady-state posterior covariance. Then, for any causal state estimator sequence $\{\psi_{t}\}_{t\geq 0}$ with $\psi_{t}\in\mathcal{F}_{t}$, we have
$\liminf_{T\to\infty}\ \frac{1}{T}\sum_{t=1}^{T}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})$ $\geq\operatorname{Tr}[\Sigma_{x,\infty}^{*}],$ (32)
almost surely. Moreover, the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}$ achieves these bounds with equality:
$\begin{split}\lim_{t\to\infty}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})&=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],\\ \lim_{T\to\infty}\ \frac{1}{T}\sum_{t=1}^{T}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})&=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],\end{split}$
almost surely. Hence, the steady-state DRKF is asymptotically minimax-optimal with respect to the worst-case one-step and long-run average MSE.
The proof of this theorem is provided in Appendix D.3. This result shows that the steady-state DRKF attains the smallest possible worst-case asymptotic MSE among all causal estimators, thereby extending the fundamental optimality of the classical KF to the DR setting. Importantly, the result concerns asymptotic performance rather than global finite-horizon minimax optimality, which is generally not well-posed under stage-wise ambiguity.
Remark 6.
@@ 4.4 Optimality
When $\theta_{v}=\theta_{w}=0$, the ambiguity sets contain only the nominal distributions. The least-favorable covariances then coincide with the nominal ones, so the DRKF reduces exactly to the classical KF. Thus, Theorem˜5 recovers the standard optimality of the KF as a special case, demonstrating that the DRKF is a strict generalization of the classical optimal estimator to the DR regime.
Overall, the results in this section confirm that the proposed noise-centric DRKF preserves the essential analytical structure of the classical KF while extending it to systems subject to distributional ambiguity. The filter remains stable, convergent, and minimax optimal, thereby bridging stochastic state estimation and DRO. These theoretical guarantees demonstrate that DRKF is not only practical but also provably reliable under distributional and long-horizon uncertainty.

### v3

@@ 4.2 Convergence
$r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d})=\alpha_{N,t}\Bigl((\Sigma_{x,t}^{-,d})^{-1}+\Omega_{N,t}\Bigr)^{-1}(\alpha_{N,t})^{\top}+W_{N,t},$ (23)
with
$\alpha_{N,t}$ $\coloneq A^{N}-\mathcal{R}_{N}\mathcal{G}_{N,t}\mathcal{O}_{N}$
$\Omega_{N,t}$ $\coloneq\mathcal{O}_{N}^{\top}(\Sigma_{u,t}^{N})^{-1}\mathcal{O}_{N}$
$W_{N,t}$ $\coloneq\mathcal{R}_{N}\mathcal{Q}_{N,t}\mathcal{R}_{N}^{\top}$
$\mathcal{Q}_{N,t}$ $\coloneq\Sigma_{w,t}^{N,*}-\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}(\Sigma_{u,t}^{N})^{-1}\mathcal{H}_{N}\Sigma_{w,t}^{N,*}$
$\mathcal{G}_{N,t}$ $\coloneq\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}(\Sigma_{u,t}^{N})^{-1}.$
Under Assumption˜2, the one-step DR Riccati update is time-invariant, and the corresponding downsampled mapping $r_{\mathbb{D},t}^{d}$ coincides with the NN-fold iterate $r_{\mathbb{D}}^{N}$ of a single Riccati operator $r_{\mathbb{D}}$. We therefore analyze contraction properties of the DR Riccati recursion through this lifted representation, via a downsampling contraction argument in the spirit of [36].
We first establish basic structural properties of $\Omega_{N,t}$ and $W_{N,t}$, which are required for the contraction analysis.
Lemma 3.
Suppose Assumptions˜3, 1 and 2 hold. Then, $W_{N,t}\in\mathbb{S}^{n_{x}}_{++}$ for any $N\geq 1$ and $\Omega_{N,t}\in\mathbb{S}^{n_{x}}_{++}$ for any $N\geq n_{x}$.
The proof is provided in Appendix C.1.
Using the above properties, we establish contraction of the (downsampled) DR Riccati mapping with respect to Thompson’s part metric, defined as follows: for $P,Q\in\mathbb{S}^{n_{x}}_{++}$,
$d_{T}(P,Q)\coloneq\log\left(\max\{\lambda_{\max}(P^{-1}Q),\lambda_{\max}(Q^{-1}P)\}\right).$
This metric is invariant under congruence transformations and is well suited for analyzing Riccati-type mappings.
Lemma 4.
Suppose Assumptions˜3, 1 and 2 hold and let $N\geq n_{x}$, so that $\mathcal{O}_{N}$ has full column rank. Then, for any $t\geq 0$, the downsampled DR Riccati mapping $r_{\mathbb{D},t}^{d}$ (23) is strictly contractive on $\mathbb{S}^{n_{x}}_{++}$ with respect to $d_{T}$: there exists $\xi_{N,t}\in(0,1)$ such that
$d_{T}(r_{\mathbb{D},t}^{d}(P),r_{\mathbb{D},t}^{d}(Q))\leq\xi_{N,t}\,d_{T}(P,Q),\quad\forall P,Q\in\mathbb{S}^{n_{x}}_{++}.$
Moreover, the contraction factor satisfies the bound
$\xi_{N,t}\leq\left(\frac{\sqrt{\|M_{N,t}\|_{2}}}{1+\sqrt{1+\|M_{N,t}\|_{2}}}\right)^{2},$ (24)
where $M_{N,t}\coloneq\Omega_{N,t}^{-1/2}\alpha_{N,t}^{\top}W_{N,t}^{-1}\alpha_{N,t}\Omega_{N,t}^{-1/2}$.
This follows from [14, Thm. 5.3] together with Lemma˜3.
We next show that the contraction factor can be bounded uniformly over time by exploiting the spectral envelope induced by Wasserstein ambiguity.
Proposition 2.
Suppose that Assumptions˜1, 2 and 3 hold. Fix $N\geq n_{x}$ and define
$\begin{split}\kappa_{N}(\theta_{v},\theta_{w})\coloneq\;&\frac{\overline{\lambda}(\hat{\Sigma}_{v},\theta_{v})+\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})\|\mathcal{H}_{N}\|_{2}^{2}}{\sigma_{\min}(\mathcal{O}_{N})^{2}}\\ &\times\left(\|A^{N}\|_{2}+\frac{\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})\|\mathcal{H}_{N}\|_{2}\|\mathcal{R}_{N}\|_{2}\|\mathcal{O}_{N}\|_{2}}{\lambda_{\min}(\hat{\Sigma}_{v})}\right)^{2}\left(\frac{1}{\lambda_{\min}(\hat{\Sigma}_{w})}+\frac{\|\mathcal{H}_{N}\|_{2}^{2}}{\lambda_{\min}(\hat{\Sigma}_{v})}\right),\end{split}$ (25)
where $\overline{\lambda}(\cdot,\cdot)$ is defined as in Proposition˜1. Then, for all $t\geq 0$, the contraction factor in (24) satisfies
$\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})\coloneq\left(\frac{\sqrt{\kappa_{N}(\theta_{v},\theta_{w})}}{1+\sqrt{1+\kappa_{N}(\theta_{v},\theta_{w})}}\right)^{2}\in(0,1).$
In particular, for any prescribed $\rho\in(0,1)$, any radii $(\theta_{v},\theta_{w})$ satisfying
$\kappa_{N}(\theta_{v},\theta_{w})\leq\frac{4\rho}{(1-\rho)^{2}}$ (26)
ensure $\xi_{N,t}\leq\rho$ for all $t$.
The proof is provided in Appendix C.2.
Combining the pointwise contraction in Lemma˜4 with the uniform bound in Proposition˜2 yields convergence of the DR Riccati recursion.
Theorem 3.
Suppose Assumptions˜3, 1 and 2 hold and let $N\geq n_{x}$. Then, for any $t\geq 0$, the downsampled DR Riccati mapping $r_{\mathbb{D},t}^{d}$ is strictly contractive on $\mathbb{S}^{n_{x}}_{++}$ with respect to Thompson’s part metric, with contraction factor $\xi_{N,t}$ satisfying
$\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})<1\qquad\forall t\geq 0.$
Consequently, the full DR Riccati recursion (18) converges to a unique fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and the associated DR Kalman gain $K_{t}$ converges.
The proof is provided in Appendix C.3.
As a consequence of Theorem˜3, the steady-state solution can be obtained by solving the single stationary SDP (13) offline, rather than via online Riccati iterations.
Corollary 3.
Under the assumptions of Theorem˜3, the steady-state covariance matrices $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$, together with the least-favorable noise covariances $\Sigma_{w,\infty}^{*}$ and $\Sigma_{v,\infty}^{*}$, solve the stationary SDP (13). Consequently, $K_{\infty}^{*}=\lim_{t\to\infty}K_{t}$.
The proof can be found in Appendix C.4.
Under Assumption˜3, the Riccati recursions of the standard KFs in Theorem˜2 converge. Taking limits in the matrix inequalities $A\preceq B$ (in the sense of the Loewner partial order) appearing in Theorem˜2, and using that the positive semidefinite cone is closed, yields the following steady-state bounds.
Corollary 4.
Under the assumptions of Theorem˜3, the steady-state prior and posterior error covariances $\Sigma_{x,\infty}^{-}$ and $\Sigma_{x,\infty}$ satisfy
$\underline{\Sigma}_{x,\infty}^{-}\preceq\Sigma_{x,\infty}^{-}\preceq\overline{\Sigma}_{x,\infty}^{-},\qquad\underline{\Sigma}_{x,\infty}\preceq\Sigma_{x,\infty}\preceq\overline{\Sigma}_{x,\infty},$
where $\underline{\Sigma}_{x,\infty}^{-}\coloneq\lim_{t\to\infty}\underline{\Sigma}_{x,t}^{-}$ and $\overline{\Sigma}_{x,\infty}^{-}\coloneq\lim_{t\to\infty}\overline{\Sigma}_{x,t}^{-}$ (and similarly for $\underline{\Sigma}_{x,\infty},\overline{\Sigma}_{x,\infty}$) are the unique steady-state limits of the lower- and upper-bound KF recursions in Theorem˜2.
@@
Since $\overline{\lambda}(\hat{X},\theta)$ is nondecreasing in $\theta$ for $\theta\geq 0$, the bound $\kappa_{N}(\theta_{v},\theta_{w})$ in (25) is monotone in both ambiguity radii. Consequently, a practical tuning procedure for $(\theta_{v},\theta_{w})$ consists of enforcing the rate constraint (26) using one of the following strategies: (i) fix $\theta_{v}$ and determine the largest admissible $\theta_{w}$; (ii) fix $\theta_{w}$ and determine the largest admissible $\theta_{v}$; or (iii) restrict $(\theta_{v},\theta_{w})$ to a one-dimensional curve (e.g., $\theta_{w}=\gamma\theta_{v}$ for some $\gamma>0$) and select the maximal radii satisfying (26). These procedures allow one to balance robustness and convergence speed in a transparent and computationally tractable manner.
4.3 Stability
A key property of the classical Kalman filter is the stability of its steady-state estimation error dynamics: under detectability of $(A,C)$ and positive definite process noise, the closed-loop error matrix (IKC)A(I-KC)A is Schur stable, ensuring bounded error covariance and asymptotic convergence of the estimation error in the unbiased case.
In the distributionally robust setting, however, the estimator is designed against least-favorable noise distributions within ambiguity sets. It is thus nontrivial whether the resulting DRKF, whose steady-state gain is characterized by the minimax design in Theorem˜3, preserves the stability guarantees of the classical KF. The following result shows that, despite the adversarial nature of the noise model, the steady-state DRKF remains asymptotically stable in the same sense as the classical KF.
@@ 4.3 Stability
Suppose Assumptions˜3, 1 and 2 hold. Let $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$ and $(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*})$ denote the steady-state DRKF covariances, and let $K_{\infty}^{*}$ be the corresponding steady-state DR Kalman gain. Define the closed-loop error matrix $F_{\infty}\coloneq(I-K_{\infty}^{*}C)A$. Then, $F_{\infty}$ is Schur stable, and the posterior estimation error $e_{t}:=x_{t}-\bar{x}_{t}$ evolves according to
[…]
Consequently, under the least-favorable noise distributions, the estimation error is mean-square stable with
$\sup_{t\geq 0}\,\mathbb{E}[\|e_{t}\|^{2}]\;<\;\infty,\quad\text{and}\quad\lim_{t\to\infty}\mathrm{cov}(e_{t})=\Sigma_{x,\infty}^{*}.$
The proof is provided in Appendix D.1.
While Theorem˜4 establishes mean-square stability, it does not by itself characterize the behavior of the estimation-error mean under model mismatch. In practice, the true noise processes may have means that differ from the nominal values used by the filter, leading to a steady-state bias. The following corollary characterizes this bias under the true noise distributions.
Corollary 5.
[…]
In particular, if the nominal noise means used by the filter coincide with the true noise means ($\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$), then $\mathbb{E}[e_{t}]\to 0$.
The proof is provided in Appendix D.2.
As shown in Lemma˜1, mean shifts are suboptimal for the adversary under Wasserstein ambiguity at the design stage, so the least-favorable noise distributions preserve the nominal means. Therefore, when the nominal noise means used by the filter coincide with the true noise means, the steady-state DRKF is unbiased in operation. When this condition is violated, a constant steady-state bias appears, exactly as characterized in Corollary˜5.
4.4 Optimality
Having established convergence and stability of the proposed DRKF, we now address its performance optimality. In the presence of distributional uncertainty, a natural question is whether one can design a causal estimator that outperforms the DRKF in terms of worst-case estimation accuracy.
In this subsection, we show that this is not possible asymptotically: among all causal estimators, the steady-state DRKF minimizes the worst-case one-step conditional mean-square error in the limit, as well as its long-run time average.
Theorem 5.
Suppose Assumptions˜1, 2 and 3 hold. Let $\Sigma_{x,\infty}^{-,*}\in\mathbb{S}^{n_{x}}_{++}$ denote the unique fixed point of the DR Riccati map $r_{\mathbb{D}}$ and let $\Sigma_{x,\infty}^{*}$ be the corresponding steady-state posterior error covariance. Then, for any causal state estimator sequence $\{\psi_{t}\}_{t\geq 0}$ with $\psi_{t}\in\mathcal{F}_{t}$, the following inequalities hold almost surely with respect to $\sigma(\mathcal{Y}_{t-1})$:
$\liminf_{T\to\infty}\ \frac{1}{T}\sum_{t=0}^{T-1}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})$ $\geq\operatorname{Tr}[\Sigma_{x,\infty}^{*}].$ (28)
Moreover, the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}^{*}$ achieves these bounds with equality:
$\lim_{t\to\infty}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})$ $=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],$ (29)
$\lim_{T\to\infty}\ \frac{1}{T}\sum_{t=0}^{T-1}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})$ $=\operatorname{Tr}[\Sigma_{x,\infty}^{*}].$ (30)
Hence, the steady-state DRKF is asymptotically minimax-optimal with respect to the worst-case one-step and long-run average MSE.
The proof of is provided in Appendix D.3. This result shows that the steady-state DRKF attains the smallest possible worst-case asymptotic MSE among all causal estimators, thereby extending the fundamental optimality of the classical KF to the DR setting. Importantly, the theorem concerns asymptotic performance: establishing finite-horizon minimax optimality under stage-wise ambiguity leads to a substantially more intricate dynamic game and is beyond the scope of this work.
@@ 4.4 Optimality
When $\theta_{v}=\theta_{w}=0$, the ambiguity sets reduce to the nominal noise distributions. In this case, the least-favorable covariances coincide with the nominal ones, and the DRKF reduces exactly to the classical KF. Thus, Theorem˜5 is consistent with the classical KF optimality and recovers it as a special case in the asymptotic regime.